OpenStudy (anonymous):

Help asap! What is the slope and y-intercept of the graph of 3x + 2y = 12? slope = negative 3 over 2; y-intercept = (0, 6) slope = -3; y-intercept = (0, 12) slope = negative 2 over 3; y-intercept = (0, 6) slope = 6; y-intercept = (0, negative 3 over 2)

OpenStudy (owlcoffee):

the ecuation is called "slope- ordenate" all you have to do is look at the ecuation and take the numbers that belong: for example: 1) m= -3/2 and interception (0.6) Use the ecuation \[y-y _{1}=m(x-x _{1})\] remplace and solve.